OpenStudy (felisha):

Can someone please help me solve for b1: A=(b1+b2/2)h

OpenStudy (raden):

it is like the formula of area of trapezium, right ?

OpenStudy (felisha):

yes

OpenStudy (raden):

well, so we have an equation : A=((b1+b2)/2)h first, multiply by 2 to both sides giving us; 2A = (b1+b2)h then divided by h to both sides : 2A/h = b1 + b2 and the last step, subtracted by b2 to both sides : 2A/h - b2 = b1

OpenStudy (felisha):

Thank You

OpenStudy (raden):

you're welcome